# SLS-BRD: A system-level approach to seeking generalised feedback Nash equilibria

### v2

@@ Abstract
This work proposes a policy learning algorithm for seeking generalised feedback Nash equilibria in $N_{P}$-players non-cooperative dynamic games. We consider linear-quadratic games with stochastic dynamics and design a best-response dynamics in which players update and communicate a parametrisation of their state-feedback policies. Our approach leverages the System Level Synthesis framework to formulate each player’s update rule as the solution to a tractable robust optimisation problem. Under certain conditions, rates of convergence to a feedback Nash equilibrium can be established. The algorithm is showcased in exemplary problems ranging from the decentralised control of unstable systems to competition in oligopolistic markets.
Index Terms:
@@
Non-cooperative games, Feedback Nash equilibrium, Best-response dynamics, System level synthesis
I Introduction
@@
Modern cyber-physical systems are often comprised of interacting subsystems operated locally by selfish decision-making agents. Ideally, agents operate these systems according to feedback policies that optimise local objectives, while satisfying global requirements and being robust to their rivals’ interference. However, the large-scale and decentralised nature of most applications, alongside the usual lack of coordination between agents, hinders most traditional approaches to policy design. Dynamic game theory provides an alternative framework based on the concept of competitive equilibria (e.g., the Nash equilibrium [1, 2]), which describe locally optimal, yet strategically stable, operating conditions for each non-cooperative agent. A methodology for policy design based on Nash equilibrium seeking thus presents a promising venue for enabling a decentralised control of cyber-physical systems.
@@
In general, solving a non-cooperative game can concern distinct goals: i) For the unbiased observer, to examine the behaviour of rational agents from a local to a global scale; ii) for the game designer, to actuate the self-interested players towards system-wide objectives; and iii) for the player, to determine locally optimal and robust policies in competitive environments. Regardless, computing a Nash equilibrium (NE) solution is a notoriously difficult task [3]. Algorithmic game theory thus emerges as the field concerned with designing procedures to bridge this computational gap [4]. An important class of algorithms are those which place to the players the task of converging towards a non-cooperative equilibrium: These include best-response dynamics (BRD, [5, 6, 7, 8, 9, 10]) and no-regret learning (NRL, [11, 12, 13, 14, 15]). Recently, online feedback optimisation has also been employed for equilibrium seeking problems [16]. In general, these are fixed-point iteration methods centred on players updating their strategies, simultaneously and independently, using only the information available to them. In particular, best-response dynamics stands out as a simple, yet fundamental, model of policy learning for uncoordinated but communicating players. These algorithms have become important tools in economics and engineering, with applications including the design and control of networked systems [17, 18, 19], robotics [20, 21], and resource management [22, 23, 24].
For dynamic games, any solution concept depends qualitatively on a description of the data available to each player; a notion termed information pattern [1]. Under open-loop information patterns, when players only access the initial state of the game, the analysis and solution of NE problems are well-understood. Conversely, games under closed-loop (or feedback) information patterns, when players monitor the state of the game, are still under active research. A feedback Nash equilibrium (FNE) solution is often more attractive than its open-loop counterpart: It renders the players robust to disturbances and decision errors. However, designing routines for computing FNE is demanding for all but simple cases. A notably challenging class of games consists of generalised feedback Nash equilibrium (GFNE, [2]) problems: When players’ policies are required to satisfy coupled constraints (e.g., restrictions on the state of the game). To the best of our knowledge, methods for FNE seeking exist only for linear-quadratic games under restrictive assumptions [25, 26, 27, 28], unconstrained nonlinear and control-affine games [29, 30, 31, 32], and application-specific problems [33, 34]. Recently, [35] expanded these results and presented a systematic (approximate) solution for GFNE problems. Despite remarkable, these efforts are restricted to finite-duration games and thus exclude decision-making processes in which agents are continuously operating the underlying system. Finally, a systematic solution is also lacking for GFNE games described by stochastic dynamics.
In this work, we investigate algorithms for GFNE seeking in (stationary) difference games with linear-stochastic dynamics. Leveraging the System Level Synthesis (SLS, [36]) framework, we propose a best-response approach based on a parametrisation of all stabilising policies for the game. The algorithm consists of players iteratively, and simultaneously, updating the parameters of their individual policies, then announcing these changes through some communication network (Figure 1). Under our approach, each player’s update rule is formulated as the solution to a tractable robust optimisation problem. Moreover, these problems allow for constraints to be imposed directly on the policies’ structure, thus having the ability to encode information patterns at the synthesis level. Finally, this proposed learning dynamics do not depend on the state and actions applied to the system, and thus can be performed simultaneously with the game’s execution. In summary, our contributions are:
@@
    A system-level best-response dynamics (SLS-BRD) algorithm for GFNE seeking in linear-stochastic dynamic games. Specifically, we design a class of best-response maps which compute an optimal system level parametrisation of a player’s policy in response to its rivals’ choices.
@@
    A realisation of the best-response mappings as robust finite-dimensional programs amenable to numerical solutions. Considering these best-responses, the SLS-BRD algorithm is formally an $\varepsilon$-GFNE seeking algorithm for an equilibrium gap $\varepsilon>0$ which can be explicitly established.
@@
    For a specific but important class of dynamic games, an analysis of the conditions and rates of convergence of the SLS-BRD to feedback Nash equilibria.
@@
This policy learning algorithm is showcased in simulated experiments on the decentralised control of an unstable network and price management in a competitive oligopolistic market.
Figure 1: SLS-BRD: Control architecture and the learning dynamics.
@@
The paper is organised as follows: Section II overviews the classes of (generalised) static and dynamic games, and the best-response dynamics algorithm. In Section III, we provide a system level parametrisation for linear-stochastic games, then design a best-response dynamics for GFNE seeking. Finally, Section IV illustrates this learning dynamics in simulated examples, and Section V provides some concluding remarks. Towards a concise presentation, only some essential theorems are proved in the main text: The remaining are in the Appendix.
I-A Notation
@@
We use Latin letters to denote vectors (lowercase) and mappings (uppercase), and use boldface to distinguish signals, operators, and their respective spaces. Sets are in calligraphic font; exceptions are the usual $\mathbb{R}$ and $\mathbb{N}$, and the sets of $N\times N$ symmetric ($\mathbb{S}^{N}$), positive semidefinite ($\mathbb{S}^{N}_{+}$), and positive definite matrices ($\mathbb{S}^{N}_{++}$). In particular, sequences are denoted $\bm{x}=(x_{t})_{t\in\mathcal{I}}$ for a countable set $\mathcal{I}\subseteq\mathbb{N}$, or $\bm{x}=(x_{t})_{t=0}^{T}$ if =$\mathcal{I}=\{0,\ldots,T\}$. For $p\in(0,\infty)$, we use the space of $N_{x}$-dimensional vector-sequences $\ell^{N_{x}}_{p}(\mathcal{I})=\{\bm{x}:\|\bm{x}\|_{\ell_{p}}=(\sum_{t\in\mathcal{I}}\|x_{t}\|^{p})^{1/p}<\infty\}$, with $\ell^{N_{x}}_{\infty}$ and ,$\ell^{N_{x}}_{\infty,e}$ the space of all bounded sequences and all sequences, respectively. $\mathcal{L}(\bm{\mathcal{X}},\bm{\mathcal{Y}})$ is the set of all bounded linear operators $\bm{A}:\bm{\mathcal{X}}\to\bm{\mathcal{Y}}$ and we sometimes denote transformed signals by $\bm{A}\bm{x}=(Ax_{t})_{t\in\mathcal{I}}$. Whenever $(\bm{\mathcal{X}},\bm{\mathcal{Y}})$ are finite-dimensional, we let $M_{A}$ refer to a matrix representation of $\bm{A}$. We use the standard definition of Hardy spaces $\mathcal{H}_{\infty}$ and $\mathcal{RH}_{\infty}$, and write $\frac{1}{z}\mathcal{RH}_{\infty}$ to refer to the set of real-rational strictly proper transfer functions. Finally, some standard signals and operators used in this paper are: The impulse signal $\bm{\delta}=(\delta_{t})_{t\in\mathcal{I}}$, the identity operator $I$ and the identity matrix $I_{N_{x}}$, and the shift operator $\bm{S_{+}}:(x_{0},x_{1},\ldots)\mapsto(0,x_{0},\ldots)$.
We distinguish set-valued mappings from ordinary functions using the notation $F:\mathcal{X}\rightrightarrows\mathcal{Y}$. For any tuple $s=(s^{p})_{p\in\mathcal{P}}\in\mathcal{S}$ we frequently write $s=(s^{p},s^{-p})$ to highlight the element $s^{p}$; this should not be interpreted as a re-ordering. Similarly, if $\mathcal{S}=\prod_{p\in\mathcal{P}}\mathcal{S}^{p}$, we define the product $\mathcal{S}^{-p}=\prod_{\tilde{p}\in\mathcal{P}\backslash\{p\}}\mathcal{S}^{\tilde{p}}$.
II Non-cooperative games and best-response dynamics
@@
defines the problem in which players $p\in\mathcal{P}=\{1,\ldots,N_{P}\}$ each decides on a strategy $s^{p}\in S^{p}(s^{-p})\subseteq\mathcal{S}^{p}$ to minimise an objective function $L^{p}:\mathcal{S}^{1}\times\cdots\times\mathcal{S}^{N_{P}}\to\mathbb{R}$. The strategy spaces $S^{p}$ ($\forall p\in\mathcal{P}$) determine the actions available to the players, with the mappings $S^{p}:\mathcal{S}^{-p}\rightrightarrows\mathcal{S}^{p}$ restricting this choice based on the actions from their opponents. As such, both the players’ objectives and feasible strategies depend explicitly on their rivals’ strategies. Finally, the players are assumed to be rational, non-cooperative, and acting simultaneously.
A solution to the game $\mathcal{G}$ is understood as a strategy profile $s=(s^{1},\ldots,s^{N_{P}})\in\mathcal{S}$, $\mathcal{S}=\mathcal{S}^{1}\times\cdots\times\mathcal{S}^{N_{P}}$, having some specified property that makes it agreeable to all players if they act rationally. In non-cooperative settings, a widely accepted solution concept is that of a generalized Nash equilibrium: The game is solved when no player can improve its objective by unilaterally deviating from the agreed strategy profile. Formally,
@@
A strategy profile $s^{*}=(s^{1^{*}},\ldots,s^{N_{P}^{*}})\in\mathcal{S}$ is a generalized Nash equilibrium (GNE) for the game $\mathcal{G}$ if
$L^{p}(s^{p^{*}},s^{-p^{*}})\leq\min_{s^{p}\in S^{p}(s^{-p^{*}})}L^{p}(s^{p},s^{-p^{*}}).$ (2)
@@
is not a singleton. As such, players might favour different solutions based on their individual objectives. We thus seek to characterise which solutions are “acceptable”, in the sense that no other equilibrium can improve the objective of all players simultaneously. Consider the partial ordering of $\Omega_{\mathcal{G}}$ defined by
$s^{*}\prec\tilde{s}^{*}\leavevmode\nobreak\ \Longleftrightarrow\leavevmode\nobreak\ L^{p}(s^{p^{*}})\leq L^{p}(\tilde{s}^{p^{*}}),\quad\forall p\in\mathcal{P},$ (3)
with at least one inequality being strict, for any pair $s^{*},\tilde{s}^{*}\in\Omega_{\mathcal{G}}$. A solution $s^{*}\in\Omega_{\mathcal{G}}$ is then characterised as an admissible GNE if there are no other equilibrium $\tilde{s}^{*}\in\Omega_{\mathcal{G}}$ such that $\tilde{s}^{*}\prec s^{*}$. Note that an admissible GNE can still be favourable to only a subset of players (that is, $\tilde{s}^{*}\nprec s^{*}\nprec\tilde{s}^{*}$), and thus it is possible that $\mathcal{G}$ has no solution that is “fair” to all players. The game might also not admit any GNE (i.e., $\Omega_{\mathcal{G}}=\emptyset$), then being characterised as unsolvable. Hereafter, we ensure that the problems being discussed are well-posed by considering the following conditions on their primitives:
@@
We investigate algorithms for solving $\mathcal{G}$. A direct computation of a GNE is equivalent to solving $N_{P}$ optimisation problems simultaneously, as implied by Definition 1. Such an approach would require players to be coordinated and their objectives to be public. Conversely, we consider adaptive procedures in which the players learn their GNE strategies independently. In this direction, consider that $\mathcal{G}$ admits episodic repetitions, and let $s_{k}\coloneqq(s_{k}^{1},\ldots,s_{k}^{N_{P}})$ be the strategy profile taken by players $\mathcal{P}$ at the $k$-th episode. A prototypical learning procedure for equilibrium seeking is outlined in Algorithm 1, where
  • $T^{p}:\mathcal{S}^{p}\times\mathcal{S}^{-p}\rightrightarrows\mathcal{S}^{p}$ describes how the $p$-th player updates its strategy, based on its opponents’ predicted next actions and given its individual objective;
  • $R^{p}:\mathcal{S}\rightrightarrows\mathcal{S}^{-p}$ describes how the $p$-th player’s predicts its opponents’ strategies for the next episode, based on the strategy profile currently being played.
@@
Output: GNE $s^{*}=(s^{1^{*}},\ldots,s^{N_{P}^{*}})$
@@
2 Initialize $s_{0}=(s_{0}^{1},\ldots,s_{0}^{N_{P}})$ and $k=0$;
3 for $k=1,2,\ldots$ do
4       if $s_{k}\in T(s_{k})$ then return $s_{k}$;

### v3

@@ Abstract
This work proposes a policy learning algorithm for seeking generalised feedback Nash equilibria (GFNE) in $N_{P}$-player noncooperative dynamic games. We consider linear-quadratic games with stochastic dynamics and design a best-response dynamics in which players update and broadcast a parametrisation of their state-feedback policies. Our approach leverages the System Level Synthesis (SLS) framework to formulate each player’s update rule as the solution to a robust optimisation problem. Under certain conditions, rates of convergence to a feedback Nash equilibrium can be established. The algorithm is showcased in exemplary problems ranging from the decentralised control of unstable systems to competition in oligopolistic markets.
This is the Accepted Manuscript of an article to appear in the IEEE Transactions on Automatic Control, available at: https://doi.org/10.1109/TAC.2025.3568560
{IEEEkeywords}
@@
Noncooperative games, best-response dynamics, feedback Nash equilibrium, system level synthesis
1 Introduction
[…]
@@ 1 Introduction
Modern cyber-physical systems are often comprised of interacting subsystems operated locally by noncooperative decision-making agents. Ideally, each agent operate its subsystem according to a feedback policy that optimise local objectives, while satisfying global constraints and being robust to their rivals’ interference. However, the large-scale, decentralised, and multi-objective nature of such applications hinders most traditional approaches to policy design. Dynamic game theory provides an alternative framework through the concept of noncooperative equilibria (e.g., the generalized Nash equilibrium [1, 2]) describing efficient, yet strategically stable, strategies for each agent. Under a dynamic game setting, decision-making agents must design feedback policies (i.e., strategies) to operate their subsystems while aware that their choice affects (and is affected by) the other agents’ choice. An equilibrium would then correspond to a set of policies that satisfy the global constraints and is agreeable to all agents given their objectives. A policy design based on generalized Nash equilibrium seeking thus presents a promising venue for the decentralised control of cyber-physical systems.
@@
In general, solving a noncooperative game has distinct goals: i) For the agents, to design efficient and robust policies in competitive environments; ii) for the game designer, to examine (and potentially control) the local and global behaviour of rational agents. Regardless, obtaining a Nash equilibrium (NE) is a notoriously difficult task [3]. Algorithmic game theory thus emerges as the field concerned with designing methods to bridge this computational gap [4]. An important class of algorithms, denoted equilibrium-seeking methods, places the task of searching for an equilibrium on the players: These include best-response dynamics (BRD, [5, 6, 7]), no-regret learning (NRL, [8, 9, 10]), and operator-splitting [11, 12] methods. Broadly, these are fixed-point methods designed for (repeated) static games centred on the idea of players improving their strategies using only the information available to them. Aside from enabling the computation of Nash equilibria, these routines have the advantage of mimicking how noncooperative players would learn their policies in reality. In particular, best-response dynamics stands out as a simple, yet fundamental, model of policy learning for uncoordinated but communicating players. This method has become an important tool in economics and engineering, with applications in networked systems [13, 14], robotics [15, 16, 17], and resource management [18, 19].
We are interested in Nash equilibrium seeking algorithms for dynamic games in which the underlying system has linear, stochastic, and potentially unstable dynamics. The focus is on closed-loop perfect state information structures, as they yield equilibrium policies that are less sensitive to modelling and decision errors than their open-loop counterparts [1]. Specifically, we consider the relevant task of players learning a generalised feedback Nash equilibrium (GFNE) of state-feedback policies which: i) stabilise the system against disturbances, ii) satisfy constraints on the control and state signals, and iii) incorporate some specific structure (e.g., encoding communication delays). In the current practice, feedback Nash equilibria is obtained by either applying dynamic programming principles [1, 20, 21, 22, 23, 24], deriving equivalent complementary problems [25, 26], or using iterative heuristics [27, 28]. Recently, [29] extended the scope to provide a systematic method to compute (approximate) GFNE. While remarkable, these solutions cannot enforce either closed-loop stability, operational constraints, or design of the policy structure, in practice. Moreover, most of them still require some central coordinator to solve the game; the players themselves do not perform equilibrium-seeking. The available literature on decentralised policy learning is concentrated on the field of reinforcement learning, where the focus is on finite-duration games described by Markov decision processes [30, 31, 32, 33, 34]. To the best of our knowledge, there are no equilibrium-seeking solutions to the aforementioned class of GFNE problems.
In this work, we propose a GFNE seeking algorithm to bridge this research gap. Leveraging the System Level Synthesis (SLS, [35]) framework, we first recast each player’s policy synthesis problem as the search for an optimal system level response to disturbances. The system level response can be used to recover a corresponding optimal feedback policy, thus serving as its parametrisation. Under this setup, each player’s best-response strategy is the solution to a tractable optimisation problem which enforces closed-loop stability, operational constraints (on the control and state signals), and structural constraints (on the policy’s parameters), already at the synthesis level. We then design a best-response routine in which players update the parametrisation of their policies, in parallel, and then announce them through some communication network (Figure 1). The algorithm does not depend on the state and actions applied to the underlying system and thus can be executed simultaneously with its operation. In summary, our contributions are:
@@
    A realisation of the associated best-response mappings in GFNE problems as robust convex optimisation problems which are amenable to numerical solutions.
@@
    A system-level best-response dynamics (SLS-BRD) algorithm for GFNE seeking in dynamic games. In this equilibrium-seeking routine, players converge to a GFNE by iteratively updating the parametrisation of their policies as the best response to their rivals’ current policies.
@@
    In the absence of shared constraints, a formal analysis of the convergence properties of the SLS-BRD algorithm.
@@
This policy learning algorithm is demonstrated in simulated experiments on the decentralised control of an unstable network and price management in a competitive oligopolistic market.
Figure 1: SLS-BRD: Control architecture and the learning dynamics. A set of players ($1,\ldots,N_{P}$) control a decentralised dynamical system by measuring its state ($\bm{x}$) and then applying actions ($\bm{u}^{1},\ldots,\bm{u}^{N_{P}}$) based on state-feedback policies ($\bm{K}^{1},\ldots,\bm{K}^{N_{P}}$) whose parametrisations ($\Phi^{1},\ldots,\Phi^{N_{P}}$) are chosen as best responses to each others’ strategies, simultaneously, and broadcasted via a communication network.
@@
The paper is organised as follows: Section 2 overviews the classes of (generalised) static and dynamic games, and the best-response dynamics algorithm. In Section 3, we provide a system level parametrisation for linear-quadratic games, then design a best-response dynamics for GFNE seeking. Finally, Section 4 illustrates this approach in exemplary problems and Section 5 provides concluding remarks. Towards a concise presentation, only essential results are given in the main text: We refer to the Supplementary Material for remaining details.
1.1 Notation
@@
We use Latin letters to denote vectors and functions, and boldface to distinguish signals, operators, and their respective spaces. Sets are in calligraphic font; exceptions are the usual $\mathbb{R}$ and $\mathbb{N}$, and the sets of $(N{\times}N)$ symmetric ($\mathbb{S}^{N}$), positive semidefinite ($\mathbb{S}^{N}_{+}$), and positive definite matrices ($\mathbb{S}^{N}_{++}$). In particular, sequences are written as $\bm{x}=(x_{t})_{t\in\mathcal{I}}$ for a countable set $\mathcal{I}\subseteq\mathbb{N}$, or $\bm{x}=(x_{t})_{t=0}^{T}$ if =$\mathcal{I}=\{0,\ldots,T\}$. For $p\in(0,\infty)$, we define the space of $N_{x}$-dimensional vector-sequences $\ell^{N_{x}}_{p}(\mathcal{I})=\{\bm{x}:\|\bm{x}\|_{\ell_{p}}=(\sum_{t\in\mathcal{I}}\|x_{t}\|^{p})^{1/p}<\infty\}$, with $\ell^{N_{x}}_{\infty}$ the space of all bounded sequences. The set $\bm{\mathcal{Y}}^{\bm{\mathcal{X}}}$ denotes all relations $\bm{A}:\bm{\mathcal{X}}\to\bm{\mathcal{Y}}$ with $\mathcal{L}(\bm{\mathcal{X}},\bm{\mathcal{Y}})\subseteq\bm{\mathcal{Y}}^{\bm{\mathcal{X}}}$ being the set of bounded linear operators. We sometimes write transformed signals as $\bm{A}\bm{x}=(Ax_{t})_{t\in\mathcal{I}}$. We use the standard definition of Hardy spaces $\mathcal{H}_{\infty}$ and $\mathcal{RH}_{\infty}$, and write $\frac{1}{z}\mathcal{RH}_{\infty}$ for the set of real-rational strictly-proper transfer functions. Finally, signals and operators used in this paper include: The impulse signal $\bm{\delta}=(\delta_{t})_{t\in\mathcal{I}}$, the identity operator $I$ and matrix $I_{N_{x}}$, the matrices $\bm{1}_{n\times m}$ and $\bm{0}_{n\times m}$ of all 1’s and 0’s, respectively, the shift operator $\bm{S_{+}}:(x_{0},x_{1},\ldots)\mapsto(0,x_{0},\ldots)$, and the Kronecker and Hadamard products $\otimes$ and $\odot$, respectively.
[…]
2 Noncooperative games and best-response dynamics
@@
defines the problem in which players $p\in\mathcal{P}=\{1,\ldots,N_{P}\}$ each decides on a strategy $s^{p}\in S^{p}(s^{-p})\subseteq\mathcal{S}^{p}$ to minimise an objective function $L^{p}:\mathcal{S}^{1}\times\cdots\times\mathcal{S}^{N_{P}}\to\mathbb{R}$. The strategy spaces $S^{p}$ ($\forall p\in\mathcal{P}$) determine the actions available to the players, with the mappings $S^{p}:\mathcal{S}^{-p}\rightrightarrows\mathcal{S}^{p}$ restricting this choice based on the actions from their opponents. As such, both the players’ objectives and feasible strategies depend explicitly on their competitors’ actions. Finally, the players are assumed to be rational, noncooperative, and acting simultaneously.
A solution to the game $\mathcal{G}$ is understood as a strategy profile $s=(s^{1},\ldots,s^{N_{P}})\in\mathcal{S}$, $\mathcal{S}=\mathcal{S}^{1}\times\cdots\times\mathcal{S}^{N_{P}}$, having some specified property that makes it agreeable to all players if they act rationally. In noncooperative settings, a widely accepted solution concept is that of a generalized Nash equilibrium: The game is solved when no player can improve its objective by unilaterally deviating from the agreed strategy profile. Formally,
@@
A strategy profile $s^{\star}=(s^{1^{\star}},\ldots,s^{N_{P}^{\star}})\in\mathcal{S}$ is a generalized Nash equilibrium (GNE) for the game $\mathcal{G}$ if
$L^{p}(s^{p^{\star}},s^{-p^{\star}})\leq\min_{s^{p}\in S^{p}(s^{-p^{\star}})}L^{p}(s^{p},s^{-p^{\star}}).$ (2)
@@
is not a singleton and can include strategy profiles that favour a specific subset of players (that is, non-admissible GNEs [1]). The game might also not admit any GNE (i.e., $\Omega_{\mathcal{G}}=\emptyset$), then being characterised as unsolvable. Hereafter, we ensure that the problems being discussed are well-posed by considering the following conditions on their primitives:
@@
Output: GNE $s^{\star}=(s^{1^{\star}},\ldots,s^{N_{P}^{\star}})$
@@
2 Initialize $s_{0}\coloneqq(s_{0}^{1},\ldots,s_{0}^{N_{P}})$ and $k\coloneqq 0$;
3 for $k=0,1,2,\ldots$ do
4       if $s_{k}\in\Omega_{\mathcal{G}}$ then return $s_{k}$;